OpenStudy (anonymous):

Can someone please help, I don't know how to do this at all... Find the derivative of f(x) = -12x^2 + 9x at x = 6.

OpenStudy (anonymous):

The equation is a standard polynomial function, so we can apply the power rule. Assuming you don't know what is it: The derivative of x^n is: \[(n)x ^{n-1}\] Where n is a number greater than 0. So applying the power rule: f'(x) = -24x + 9 For the power rule, we can ignore constants, so we essentially do nothing with the -12 and leave it as it is. Then we plug in 6 into f'(x).

OpenStudy (anonymous):

Since I don't think I covered constants and the power enough: -12x^2's derivative is equal to -12 * the derivative of x^2 = -12 * 2x = -24x Doesn't matter if it's -12, 50, 100, 20, etc. We can do this for any integer.

OpenStudy (anonymous):

Alright, so the we have... -24x+9x -24(6)+9(6) -90

OpenStudy (anonymous):

I got -90 also, so confirming you're right.

OpenStudy (anonymous):

No wait, you put the wrong derivative. It's -135, because the derivative of 9x is 9, not 9x.

OpenStudy (anonymous):

Now, I'm just confused...

OpenStudy (anonymous):

A derivative is the slope of an equation. So, for 9x, if we draw the graph, we can see that it's a straight line with a slope of 9.

OpenStudy (anonymous):

The slope is constant at all pts on the line, so we put 9 as the derivative of 9x.

OpenStudy (anonymous):

OhKay then...